iamchris08:

Given the equation square root of 8x plus 1 = 5, solve for x and identify if it is an extraneous solution.

AngeI:

8x+1=5 You would have to subtract 1 from both sides, getting 8x=4 Then you would divide both sides by 8 Which means X=?

AngeI:

Welcome to QuestionCove cx

iamchris08:

whats the answer?

AngeI:

I can't directly give you the answer because then you won't learn To find the value of x all that is left for you to do is divide 4 by 8

iamchris08:

it cant be two

AngeI:

No, because 8 divided by 4 is 2 You are supposed to do 4 divided by 8

iamchris08:

x = one fourth, solution is not extraneous x = one fourth, solution is extraneous x = 3, solution is not extraneous x = 3, solution is extraneous

iamchris08:

it has to be one of these

Nnesha:

angel it's the square of "8x+1"

Nnesha:

\[\sqrt{8x+1}=5\] like this :=))

Nnesha:

square root ** of "8x+1"

AngeI:

OH

AngeI:

My apologies @iamchris08

iamchris08:

no youre good

AngeI:

Nnesha you wanna take this one?

Nnesha:

you can continue..

AngeI:

You would square both sides, getting 8x+1=25 Then subtract the 1 from both sides getting 8x=24 So x=?

iamchris08:

3

AngeI:

Then do (Square root) 8*3+1=5 Simplifying it to get 5=5 showing thay 3 is your answer c:

iamchris08:

thx